OpenStudy (anncherry):

How do you solve 3(x-k)/w = 4

KjSaif (kjsaif):

NO DIRECT ANSWERS

KjSaif (kjsaif):

:)

OpenStudy (simontheengineer):

Depends on the values of other numbers. First expand the brackets to get (3x-3k)/w is 4. You then need a value for x, k or w to help.

OpenStudy (anncherry):

solving for x

OpenStudy (dallascowboys88):

You Have To Use The Formula: PEMDAS=(P)Parenthesis,(E)Equalaties,(M)Multiply,(D)Divide,(A)Add,(S)Subtract.Exactly In That Order To Get Your Answer.

KjSaif (kjsaif):

3(x−k)/w=4 Step 1: Multiply both sides by w. −3k+3x=4w Step 2: Add -3x to both sides. −3k+3x+−3x=4w+−3x −3k=4w−3x Step 3: Divide both sides by -3. −3k/−3 = 4w−3x/−3k= −4/3 w+x

KjSaif (kjsaif):

\[k = -4\div3 w + x\]

KjSaif (kjsaif):

@anncherry